OpenStudy (anonymous):

Solve s= P/2 (c+K) for K

OpenStudy (ash2326):

\[ s= \frac{p}{2} (c+k)\] multiply both sides by 2 and divide both sides by p \[\frac{2}{p} s= c+k\] subtract both sides by k \[\frac{2}{p} s-c= k\] so k is \(\frac{2}{p} s-c\)

OpenStudy (anonymous):

Thank-you :D

OpenStudy (anonymous):

Solve x-6y=19 for y?

Directrix (directrix):

Solve x-6y=19 for y x-6y = 19 -6y = 19 - x -1/6(-6y) = (-1/6)(19 - x) y = -19/6 + x/6 y = (1/6) x - 19/6